Osmium has a density of 22.6 g/mL. Calculate the radius of an Os atom expected if osmium crystallizes in the face-centered cubic structure.


A.701 pm


B. 101 pm


C. 338 pm


D. 135 pm


1
Expert's answer
2021-10-21T02:07:50-0400

In a face-centered cubic unit cell there is total 4 atoms per cell. (8*1/8 + 2*1/2 + 2*1/2 + 2*1/2 = 4)


In 22.6 grams of osmium there is 22.6/190.23 = 0.1188 mol or 0.1188*6.02*10^23 = 7.152*10^22 atoms.


This means that 1 cm3 of osmium has 7.152*10^22 atoms.


We need to find the volume that contains 4 atoms.

V = 4/(7.152*10^22) = 5.59*10^-23 cm3 = 5.59*10^7 pm3. This is the volume of out face-centered cubic unit cell.


It is clear from the picture, that the radius of osmium atom is 1/4 of the cube diagonal. The cube edge is "\sqrt[3]{5.59*10^7}" = 382 pm. The diagonal is "\sqrt{2}*382" = 540.8. 1/4 of the diagonal is 540.8/4 = 135.2 pm.


Answer: 135 pm.